[figure: 9.7 cm]

THE ELECTRONIC STRUCTURE OF THE RMn_{2}Ge_{2} (R = Ca, Y, Ba, La) ANTIFERROMAGNETS

aJ. Toboła, ^{b}B. Malaman and ^{b}G. Venturini

aFaculty of Physics and Nuclear Techniques, Academy of Mining and Metallurgy,

Al. Mickiewicza 30, 30-059 Kraków, Poland bLaboratoire de la Chimie du Solide Mineral, Université Henri Poincaré, 54506 Vandoeuvre les Nancy Cedex, France

We present the electronic structure of the tetragonal RMn_{2}Ge_{2} (R=Ca,Y,Ba,La) antiferromagnets investigated by the selfconsistent KKR calculations. In the above-mentioned compounds the two types of antiferromagnetic structure (AF_{1} and AF_{2}) were experimentally detected. In YMn_{2}Ge_{2} the Mn magnetic moments are arranged parallel in the a-b plane, while antiferromagnetically inter the a-b planes (AF_{1}). On the contrary, in CaMn_{2}Ge_{2}, LaMn_{2}Ge_{2} and BaMn_{2}Ge_{2}, the neighboring μ_{Mn} (in the a-b plane) are aligned anti-parallel (AF_{2}). Moreover, the ordering of the Mn magnetic moment in RMn_{2}Ge_{2} seems to be governed by the Mn-Mn distance d_{in} in the a-b plane. For d_{in} < 2.87 Åthe AF_{1}-type of magnetic arrangement is observed, otherwise (d_{in} ≥ 2.87Å) the AF_{2}-type occurs. The KKR results of the RMn_{2}Ge_{2} compounds show that the μ_{Mn} value increases more or less linearly with d_{in}. The calculated magnetic moments on Mn: 2.17μ_{B} (YMn_{2}Ge_{2}), 2.74μ_{B} (CaMn_{2}Ge_{2}), 2.95μ_{B} (LaMn_{2}Ge_{2}) and 3.47μ_{B} (BaMn_{2}Ge_{2}) remain in good agreement with the neutron diffraction measurements (in μ_{B}): 2.20, 2.67, 3.05, 3.66, respectively. The reason for the changing of the AF-type structure, near d_{in} = 2.87Å, is discussed in a view of the total energy KKR computations.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Janusz Toboła

Address for correspondence : Faculty of Physics and Nuclear Techniques, AGH 30-059 Kraków, Poland

Email address : tobola@ftj.agh.edu.pl tobola@grxtal.polycnrs-gre.fr